Suppose that the scores on a reading ability test are normally distributed with a mean of

65

and a standard deviation of

10

. What proportion of individuals score at least

55

points on this test? Round your answer to at least four decimal places.
Answer #1

Solution :

Given that,

mean = = 65

standard deviation = = 10

P(x >55 ) = 1 - P(x<55 )

= 1 - P[(x -) / < (55-65) / 10]

= 1 - P(z <-1 )

Using z table

= 1 - 0.1587

probability=08413